you receive a total of $1520 a year in interest from three investments. the interest rate for the three investments are 5%, 7%, and 8%. the 5% investment is half of the 7% investment, and the 7% investment is $1500 less than the 8%. What is the amount of each investment?

(Note: This is a mixture problem. So set up a table involving Amount (of investment), Rate, and Interest. Let us use Investment A for 5%, Investment B for 7%, and Investment C for 8%)

Answer by lwsshak3(11628)

you receive a total of $1520 a year in interest from three investments. the interest rate for the three investments are 5%, 7%, and 8%. the 5% investment is half of the 7% investment, and the 7% investment is $1500 less than the 8%. What is the amount of each investment?
***
let A=amt of 5% investment
let B=amt of 7% investment
let C=amt of 8% investment
..
B=C-1500
A=(C-1500)/2
..
5%*(C-1500)/2+7%(C-1500)+8%C=1520
lcd:2
5%(C-1500)+14%(C-1500)+16%C=3040
.05C-75+.14C-210+.16C=3040
.35C=3325
C=9500
B=C-1500=8000
A=B/2=4000
amt of 5% investment=$4000
amt of 7% investment=$8000
amt of 8% investment=$9500